# Learning from Negative Samples in Biomedical Generative Entity Linking

### v2

@@ Abstract
Chanhwi Kim1*, Hyunjae Kim1*, Sihyeon Park1, Jiwoo Lee1, Mujeen Sung2, Jaewoo Kang1,3 1Korea University, 2Kyung Hee University, 3AIGEN Sciences {chanhwi_kim, hyunjae-kim, sh10, hijiwoo7}@korea.ac.kr mujeensung@khu.ac.kr, kangj@korea.ac.kr
@@ 2.1 Biomedical Entity Linking
Traditional classification-based approaches Limsopatham and Collier (2016a); Miftahutdinov et al. (2019) employed a softmax layer for classification, treating concepts as categorical variables and thereby losing the detailed information of concept names. Similarity-based Sung et al. (2020); Liu et al. (2021); Lai et al. (2021); Zhang et al. (2022) models have significantly improved BioEL performance, which encodes mentions and candidate entity names in the same vector space. They are characterized by high memory consumption due to the need to encode entities into pre-computed embeddings, posing scalability challenges with large datasets De Cao et al. (2020). Several studies have integrated the concept of clustering into BioEL Angell et al. (2021); Agarwal et al. (2022).
@@ 3.2 ANGEL Framework
Our framework supports not only fine-tuning with labeled datasets but also pre-training with the KB. Specifically, we automatically generate surrounding contextual information for each entity in the KB, using clause templates or definitions, as outlined in GenBioEL Yuan et al. (2022b). When the entity definition corresponding to the identifier is available, a synonym and its definition are integrated into a pre-defined clause template as follows:222Refer to Yuan et al. (2022b) for the full set of templates.
@@ 4.1 Datasets
4.2 Baseline Models
We used top-performing similarity-based models Sung et al. (2020); Liu et al. (2021); Lai et al. (2021); Zhang et al. (2022) as our baselines. Notably, Prompt-BioEL Xu et al. (2023) employs a re-ranking-based approach. In the first stage, a similarity-based model, such as SapBERT, retrieves the top-k candidate entities from the knowledge base. In the second stage, these candidates are reranked using a cross-encoder. Although Prompt-BioEL may not be directly comparable, as it incorporates additional modules on top of existing models, we report its performance alongside for reference. Additionally, we include the previously best-performing generative models for comparison. (1) BART-large Lewis et al. (2020) is an encoder-decoder language model pre-trained on a general-domain corpus. (2) BioBART-large Yuan et al. (2022a) is the BART-large model continuously pre-trained on a biomedical-domain corpus. (3) GenBioEL Yuan et al. (2022b) is initialized with the weights of the BART-large model and then pre-trained specifically for BioEL using UMLS. We excluded several models Agarwal et al. (2022); Lin et al. (2024) due to the lack of publicly available code or the difficulty in reproducing their reported performance.
@@ 4.3 Implementation Details
Our framework was applied to each of these models during fine-tuning, referred to as ANGEL, and during both pre-training and fine-tuning, referred to as ANGEL. For pre-training, we utilized the 2020AA version of the UMLS database,333https://www.nlm.nih.gov/research/umls/licensedcontent/umlsarchives04.html which comprises 3.09M entities, of which 199K concepts contain definitions. During pre-training, we saved checkpoints every 500 steps over 5 epochs and selected the best one based on the validation sets. We used top-3 synonyms as positive samples in positive-only training. The other hyperparameter configurations are detailed in Appendix B. In pre-processing, following Yuan et al. (2022b), we expanded abbreviations using AB3P Sohn et al. (2008), lowercase texts, mark mention boundaries with special tokens [ST] and [ET], and discard mentions that overlap or are missing from the target KB. During pre-training, our models were trained using eight 80G A100 GPUs for 12 hours. During fine-tuning, a single A100 GPU was used.
@@ 5.1 Ablation Study
Similarity models play a critical role in retrieving synonymous terms from the KB, and their choice can have a substantial impact on overall system performance. To assess their effectiveness, we evaluated three models: (1) TF-IDF, (2) BioBERT-NLI,444https://huggingface.co/gsarti/biobert-nli a sentence embedding model fine-tuned on natural language inference datasets, and (3) SapBERT. These models were incorporated into the positive-aware training framework of GenBioEL on the NCBI-Disease and BC5CDR datasets. As presented in Table 3, the TF-IDF-based approach outperforms the two embedding-based models. Although the strong performance of trigram-based similarity highlights the utility of surface-level matching in BioEL, this does not imply that the task is inherently simple. While many synonyms exhibit similar surface forms, a substantial portion do not—posing challenging edge cases that demand more nuanced semantic understanding.
@@ Appendix B Hyperparameter Configurations
Table B details the hyperparameters used for positive-only training and negative-aware training across the BioEL benchmark datasets. We searched for the optimal hyperparameter settings using the validation sets. We refer to the study of Yuan et al. (2022b) to determine the range of the hyperparameters. During pre-training, we used the same hyperparameters as in GenBioEL. For positive-only training, we explored a range of training steps between 20K and 40K, a learning rate between 2e-5 and 3e-7, and batch sizes from 8 to 16, except during pre-training. During negative-aware training, we fixed the at 0.1, in accordance with the basic configuration of DPO, and searched the hyperparameter space using a learning rate between 2e-5 and 1e-6 and batch sizes ranging from 8 to 64. We used the source codes provided by Yuan et al. (2022b)555https://github.com/Yuanhy1997/GenBioEL and alignment handbook Tunstall et al. (2023)666https://github.com/huggingface/alignment-handbook.

### v3

@@ Abstract
Chanhwi Kim1*, Hyunjae Kim1*, Sihyeon Park1, Jiwoo Lee1, Mujeen Sung2, Jaewoo Kang1,3 1Korea University, 2Kyung Hee University, 3AIGEN Sciences {chanhwi_kim, hyunjae-kim, sihyeon-park, hijiwoo7}@korea.ac.kr mujeensung@khu.ac.kr, kangj@korea.ac.kr
@@ 2.1 Biomedical Entity Linking
Traditional classification-based approaches Limsopatham and Collier (2016a); Miftahutdinov et al. (2019) employed a softmax layer for classification, treating concepts as categorical variables and thereby losing the detailed information of concept names. Similarity-based Sung et al. (2020); Liu et al. (2021); Lai et al. (2021); Zhang et al. (2022) models have significantly improved BioEL performance, which encodes mentions and candidate entity names in the same vector space. They are characterized by high memory consumption due to the need to encode entities into pre-computed embeddings, posing scalability challenges with large datasets De Cao et al. (2020). integrated the concept of clustering into BioEL Angell et al. (2021); Agarwal et al. (2022).
@@ 3.2 ANGEL Framework
Our framework supports not only fine-tuning with labeled datasets but also pre-training with the KB. Specifically, we automatically generate surrounding contextual information for each entity in the KB, using clause templates or definitions, as outlined in GenBioEL Yuan et al. (2022b). When the entity definition corresponding to the identifier is available, a synonym and its definition are integrated into a pre-defined clause template as follows:111Refer to Yuan et al. (2022b) for the full set of templates.
@@ 4.1 Datasets
similarity-based models Sung et al. (2020); Liu et al. (2021); Lai et al. (2021); Zhang et al. (2022) as our baselines. Notably, Prompt-BioEL Xu et al. (2023) employs a re-ranking-based approach. In the first stage, a similarity-based model, such as SapBERT, retrieves the top-k candidate entities from the knowledge base. In the second stage, these candidates are reranked using a cross-encoder. Although Prompt-BioEL may not be directly comparable, as it incorporates additional modules on top of existing models, we report its performance alongside for reference. Additionally, we include the previously best-performing generative models for comparison. (1) BART-large Lewis et al. (2020) is an encoder-decoder language model pre-trained on a general-domain corpus. (2) BioBART-large Yuan et al. (2022a) is the BART-large model continuously pre-trained on a biomedical-domain corpus. (3) GenBioEL Yuan et al. (2022b) is initialized with the weights of the BART-large model and then pre-trained specifically for BioEL using UMLS. We excluded several models Agarwal et al. (2022); Lin et al. (2024) due to the lack of publicly available code or the difficulty in reproducing their reported performance.
@@ 4.2 Implementation Details
Our framework was applied to each of these models during fine-tuning, referred to as ANGEL, and during both pre-training and fine-tuning, referred to as ANGEL. For pre-training, we utilized the 2020AA version of the UMLS database,222https://www.nlm.nih.gov/research/umls/licensedcontent/umlsarchives04.html which comprises 3.09M entities, of which 199K concepts contain definitions. During pre-training, we saved checkpoints every 500 steps over 5 epochs and selected the best one based on the validation sets. We used top-3 synonyms as positive samples in positive-only training. The other hyperparameter configurations are detailed in Appendix B. In pre-processing, following Yuan et al. (2022b), we expanded abbreviations using AB3P Sohn et al. (2008), lowercase texts, mark mention boundaries with special tokens [ST] and [ET], and discard mentions that overlap or are missing from the target KB.g pre-training, our models were trained using eight 80G A100 GPUs for 12 hours. During fine-tuning, a single A100 GPU was used.
@@ 5.1 Ablation Study
Similarity models play a critical role in retrieving synonymous terms from the KB, and their choice can have a substantial impact on overall system performance. To assess their effectiveness, we evaluated three models: (1) TF-IDF, (2) BioBERT-NLI,333https://huggingface.co/gsarti/biobert-nli a sentence embedding model fine-tuned on natural language inference datasets, and (3) SapBERT. These models were incorporated into the positive-aware training framework of GenBioEL on the NCBI-Disease and BC5CDR datasets. As presented in Table 3, the TF-IDF-based approach outperforms the two embedding-based models. Although the strong performance of trigram-based similarity highlights the utility of surface-level matching in BioEL, this does not imply that the task is inherently simple. While many synonyms exhibit similar surface forms, a substantial portion do not—posing challenging edge cases that demand more nuanced semantic understanding.
@@ Appendix B Hyperparameter Configurations
Table B details the hyperparameters used for positive-only training and negative-aware training across the BioEL benchmark datasets. We searched for the optimal hyperparameter settings using the validation sets. We refer to the study of Yuan et al. (2022b) to determine the range of the hyperparameters. During pre-training, we used the same hyperparameters as in GenBioEL. For positive-only training, we explored a range of training steps between 20K and 40K, a learning rate between 2e-5 and 3e-7, and batch sizes from 8 to 16, except during pre-training. During negative-aware training, we fixed the at 0.1, in accordance with the basic configuration of DPO, and searched the hyperparameter space using a learning rate between 2e-5 and 1e-6 and batch sizes ranging from 8 to 64. We used the source codes provided by Yuan et al. (2022b)444https://github.com/Yuanhy1997/GenBioEL and alignment handbook Tunstall et al. (2023)555https://github.com/huggingface/alignment-handbook.